OpenStudy (anonymous):

simplify sqrt{25(t-4)^6}

OpenStudy (anonymous):

5(t-4)^3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Do you know how to do it ?

OpenStudy (anonymous):

i know how to

OpenStudy (anonymous):

No thats what I need help, I am really bad at algebra

OpenStudy (anonymous):

i can explain it to u :)

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

\[\sqrt{a}=a ^{1/2}\] \[\sqrt[3]{a}=a ^{1/3}\] I think you should be able to solve with this.

OpenStudy (anonymous):

ok